Calculate the conditional formation constant of the EDTA complex of Al3+ in a solution buffered to pH = 3.0 Answer to 3 significant digits.
Answer #1

Consider complex formation reaction, Al 3+ + Y 4-\rightleftharpoons [ Al Y -1]

The equilibrium constant for above reaction is called formation constant and is given as,

K f = [ Al Y -1] / [Al 3+ ] [Y 4-]

The conditional formation constant is given as, K ' f = \alphaY 4- K f

At pH = 3.0 , \alphaY 4- = 2.1 \times 10 -11

From literature, log K f = 16.4

\therefore Conditional formation constant ( K ' f ) = ( 2.1 \times 10 -11 ) 10 16.4 = 5.27 \times 10 5

ANSWER : K ' f = 5.27 \times 10 5